OpenStudy (rational):

\[\large (p-1)! + 1 = p^k\] Find all solutions such that \(p\) is prime and \(k\in \mathbb{Z^{+}}\)

OpenStudy (ikram002p):

so we can rearrange it like this:- \(\Large (p-2)!=\frac{p^k-1}{p-1}\)

OpenStudy (ikram002p):

any hint ?

OpenStudy (rational):

previous problem

OpenStudy (rational):

btw thats a good start!

OpenStudy (ikram002p):

seems like my way to solve pre goes to no where xD

OpenStudy (rational):

thats actually is the first step in the solution i have

OpenStudy (ikram002p):

hmm (p-1)|(p-2)!

OpenStudy (rational):

\[(2-1) \nmid (2-2)!\] \[(5-1) \nmid (5-2)!\]

OpenStudy (ikram002p):

yeah for p>5 hmm

OpenStudy (rational):

Okay for p>5 we have (p-1) | (p-2)! looks good, keep going...

OpenStudy (kainui):

I can't escape the geometric series no matter what hahaha.

OpenStudy (ikram002p):

-.- eh lol

OpenStudy (rational):

Yes! writing it as geometric series makes the next step clear

OpenStudy (kainui):

Based on what we came across the other day and using "p" \[\Large \sigma(p^{k-1})= (p-2)!\]

OpenStudy (rational):

yes this form might give some hint about next step \[\Large (p-2)!=\sum\limits_{i=0}^{k-1} p^i \]

OpenStudy (ikram002p):

hmm i would say no solution :|

OpenStudy (rational):

thats right! there are no solutions for p>5 how did you arrive at that ?

OpenStudy (ikram002p):

just a conjecture xD

OpenStudy (rational):

you're almost there... try using the result of previous problem

OpenStudy (rational):

previous prblem http://openstudy.com/users/rational#/updates/55096fc6e4b0b8a80199adcb

OpenStudy (kainui):

Because of the factorial we know that for the statement to be true this statement must also be true \[\Large 2^{\frac{p-1}{2}} | \ p^k-1\]

OpenStudy (ikram002p):

\(\Large (p-2)!=\sum\limits_{i=0}^{k-1} p^i =0 \mod (p-1) \)

OpenStudy (rational):

BINGO!

OpenStudy (rational):

kai Is that because there are (p-1)/2 even factors on left hand side factorial ?

OpenStudy (ikram002p):

\( \Large \sum\limits_{i=0}^{k-1} p^i= \sum\limits_{i=0}^{k-1} 1^i \mod (p-1)\\ \Large =1+1+...+1 \mod (p-1) = 0\mod (p-1) \) hmmmm note we have k-1 1's

OpenStudy (ikram002p):

ugh we have k 1's :2

OpenStudy (rational):

\[\sum\limits_{i=\color{Red}{0}}^{k-1}1 = k\]

OpenStudy (ikram002p):

k=0 mod p-1 k= n(p-1)

OpenStudy (kainui):

Yes exactly. =) Now since primes greater than 2 are odd, we can write them as this: \[\Large p^k-1 = (2n+1)^k-1 = \sum_{i=1}^k \binom{k}{i}(2n)^i\] however for this to be divisible by \[\Large 2^{\frac{p-1}{2}}\] then we can't have the exponent on 2 be larger than 1 since the i=1 term of our sum only has a single 2 in it.... Although now I realize this is going to be an incomplete argument since n itself could be even. However for odd n this argument holds so we have to take this a litte further I guess for this way I'm cooking up. hmm... I'll check out ikram's stuff and come back to this in a minute.

OpenStudy (ikram002p):

typo @rational :2

OpenStudy (kainui):

An extra thing I should have realized earlier, a simple limit to put on k is this, \[\Large p^p > p! \implies p^p> p^{k+1}-p\] \[\Large p^{p-1}+1>p^k\] Hmmm... just some ideas.

OpenStudy (ikram002p):

hehehe xD

OpenStudy (rational):

and ikram got k = n(p-1) we're in a good position to conclude i guess XD

OpenStudy (ikram002p):

ok so p^p>p! p^(p-1)>(p-1)! p^k>(p-1)!

OpenStudy (rational):

\[(p-1)! + 1= p^{\color{blue}{k}}\] \[(p-1)! + 1= p^{\color{blue}{n(p-1)}}\] however we have \[(p-1)! \lt p\cdot p\cdots (\text{p-1 times}) = p^{p-1}\] so there are no solutions for \(\p \gt 5\) in as much as there are no solutions for \(n \ge 1\)

OpenStudy (ikram002p):

nice :3 thanks for the question

OpenStudy (kainui):

Nice!